# MT2ST: Adaptive Multi-Task to Single-Task Learning

### v5

@@ 1 Introduction
The rapid evolution of large-scale models in machine learning (ML), particularly in natural language processing (NLP), computer vision (CV), and speech recognition, has brought tremendous advances in task performance but also increased the demand for computational efficiency. As models grow in parameter size and data requirements, efficient training strategies have become indispensable for scalable deployment and practical adaptation. Among these, the training of task-specific embeddings remains a fundamental component, serving as the backbone for semantic representation in both unimodal and multimodal applications Mikolov et al. (2013); Zhang et al. (2023a).
A major trade-off emerges in the choice of training paradigm: single-task learning (STL) vs. multi-task learning (MTL). STL enables high-fidelity adaptation to a specific task objective, often yielding superior precision. However, it lacks inductive bias and representation reuse, limiting generalization. In contrast, MTL introduces auxiliary tasks that can guide shared representation learning, promoting robustness and faster convergence, especially in low-resource regimes Collobert and Weston (2011). Nevertheless, MTL is not without cost: task interference, gradient conflict Sener and Koltun (2018), and heterogeneous learning dynamics can degrade both convergence speed and final task performance Zhang et al. (2023a); Yu et al. (2020).
@@ 2.1 Challenges in Single-Task Representation Learning
Effective representation learning typically demands large-scale datasets to capture contextual and task-relevant patterns. As datasets grow in size and complexity, training time and resource requirements increase significantly Ebner et al. (2019); Liu et al. (2024a). This presents a practical barrier to deploying scalable machine learning solutions, particularly for real-time or resource-constrained environments.
@@
Learning expressive representations often involves deep architectures and iterative optimization over millions or billions of parameters. This leads to high computational costs and energy consumption Liu et al. (2024b), prompting the need for efficient training strategies and algorithmic improvements.
@@
The optimization landscape of representation learning is typically non-convex and high-dimensional, making convergence difficult and sensitive to initialization, batch composition, and training dynamics Ganea et al. (2021); Smith et al. (2021). These challenges are amplified in real-world settings where data is noisy, multi-modal, or weakly labeled.
@@ 2.2 Improving Training Efficiency via Multi-Task Learning
Multi-task learning (MTL) is a widely adopted paradigm aimed at improving model efficiency and generalization by jointly training on multiple related tasks. In MTL, shared representations are learned across tasks, allowing the model to benefit from auxiliary supervision and mutual inductive bias Caruana (1997). MTL has proven effective across domains, including NLP Zhang et al. (2023b); Su et al. (2022), computer vision Misra et al. (2016); Zhang et al. (2020), and speech recognition.
@@ 2.3 Limitations of MTL for General Representation Learning
MTL may incur additional computational cost due to task-specific heads, losses, and gradient computations. As the number of tasks increases, these costs accumulate, reducing the practical efficiency gains of MTL Liu et al. (2020).
@@
Scaling MTL to many tasks often results in task imbalance and dominance by easier or higher-resource tasks. This imbalance can distort the shared representations and lead to underperformance on the primary task Ruder (2017); Du et al. (2021).

### v6

@@ 1 Introduction
The rapid evolution of large-scale models in machine learning (ML), particularly in natural language processing (NLP), computer vision (CV), and speech recognition, has brought tremendous advances in task performance but also increased the demand for computational efficiency. As models grow in parameter size and data requirements, efficient training strategies have become indispensable for scalable deployment and practical adaptation. Among these, the training of task-specific embeddings remains a fundamental component, serving as the backbone for semantic representation in both unimodal and multimodal applications Mikolov et al. (2013); Zhang and Yang (2021).
A major trade-off emerges in the choice of training paradigm: single-task learning (STL) vs. multi-task learning (MTL). STL enables high-fidelity adaptation to a specific task objective, often yielding superior precision. However, it lacks inductive bias and representation reuse, limiting generalization. In contrast, MTL introduces auxiliary tasks that can guide shared representation learning, promoting robustness and faster convergence, especially in low-resource regimes Wang et al. (2020); Chung et al. (2022). Nevertheless, MTL is not without cost: task interference, gradient conflict Sener and Koltun (2018), and heterogeneous learning dynamics can degrade both convergence speed and final task performance Zhang et al. (2023); Zhang and Yang (2021); Yu et al. (2020).
@@ 2.1 Challenges in Single-Task Representation Learning
Effective representation learning typically demands large-scale datasets to capture contextual and task-relevant patterns. As datasets grow in size and complexity, training time and resource requirements increase significantly Ebner et al. (2019); Liu and Pister (2024). This presents a practical barrier to deploying scalable machine learning solutions, particularly for real-time or resource-constrained environments.
@@
Learning expressive representations often involves deep architectures and iterative optimization over millions or billions of parameters. This leads to high computational costs and energy consumption Liu et al. (2024), prompting the need for efficient training strategies and algorithmic improvements.
@@
The optimization landscape of representation learning is typically non-convex and high-dimensional, making convergence difficult and sensitive to initialization, batch composition, and training dynamics Zeng and Nie (2021); Ban and Ji (2024); Zhao et al. (2023). These challenges are amplified in real-world settings where data is noisy, multi-modal, or weakly labeled.
@@ 2.2 Improving Training Efficiency via Multi-Task Learning
Multi-task learning (MTL) is a widely adopted paradigm aimed at improving model efficiency and generalization by jointly training on multiple related tasks. In MTL, shared representations are learned across tasks, allowing the model to benefit from auxiliary supervision and mutual inductive bias Caruana (1997). MTL has proven effective across domains, including NLP Zhang et al. (2023); Su et al. (2022), computer vision Lopes et al. (2024); Zhang and Yang (2021), and speech recognition.
@@ 2.3 Limitations of MTL for General Representation Learning
MTL may incur additional computational cost due to task-specific heads, losses, and gradient computations. As the number of tasks increases, these costs accumulate, reducing the practical efficiency gains of MTL Zhang et al. (2023).
@@
Scaling MTL to many tasks often results in task imbalance and dominance by easier or higher-resource tasks. This imbalance can distort the shared representations and lead to underperformance on the primary task Ruder (2017); Ahmad et al. (2018); Trabelsi et al. (2021).